3. Pyruvic acid, HC3H303, is involved in cell metabolism. It can be measured by using a yeast enzyme to complete the following reaction HC3H3O3 (aq) → C3H40 (aq) + CO2 (g) If a sample of pyruvic acid yields 25.0 ml of carbon dioxide gas at 350 mmHg and 25 °c, how many grams of pyruvic acid are in the sample? Note: at STP 1 mnolega. 22.4 L
Answer #1

1st calculate moles of CO2 formed

we have:

P = 350.0 mm Hg

= (350.0/760) atm

= 0.4605 atm

V = 25.0 mL

= (25.0/1000) L

= 0.025 L

T = 25.0 oC

= (25.0+273) K

= 298 K

find number of moles using:

P * V = n*R*T

0.4605 atm * 0.025 L = n * 0.08206 atm.L/mol.K * 298 K

n = 4.706*10^-4 mol

from reaction,

moles of HC3H3O3 reacting = moles of CO2 formed

= 4.706*10^-4 mol

Molar mass of HC3H3O3 = 3*MM(H) + 3*MM(C) + 3*MM(O)

= 3*1.008 + 3*12.01 + 3*16.0

= 87.054 g/mol

we have below equation to be used:

mass of HC3H3O3,

m = number of mol * molar mass

= 4.706*10^-4 mol * 87.054 g/mol

= 4.10*10^-2 g

Answer: 4.10*10^-2 g

Feel free to comment below if you have any doubts or if this answer do not work